what does x equal in the equation: 2-3(2x-1)=17
Answer by nyc_function(2741)

2-3(2x-1)= 17

2 -6x + 3 = 17

-6x + 5 = 17

-6x = 17 - 5

-6x = 12

x = 12/-6

x = -2

mrfeliz.webs.com